# PriPHiT: Privacy-Preserving Hierarchical Training of Deep Neural Networks

### v1

@@ VIII White-Box Reconstruction Attack
IX Privacy Budget Analysis
Figs. 13 and 14 show the effect of the privacy budget $\epsilon$ on the accuracy of the cloud analyzer in comparison to the accuracy of the cloud classification attacker for the smiling versus gender and having an open mouth versus wearing makeup experiments on the CelebA dataset, respectively. In the PriPHiT method, reducing the privacy budget decreases the classification attacker’s accuracy on the sensitive content significantly while causing a minimal cost in the analyzer accuracy on the desired content.
X Qualitative Results of the Reconstruction Attack in Mouth Open vs. Wearing Makeup Experiment
Figs. 15 and 16 show examples of the results of the reconstruction attacks together with the unseen inputs from the testing subset for the mouth open versus wearing heavy makeup experiment on the CelebA dataset. We see that the output of both attacks can reconstruct the inputs with makeup indicators using the extracted feature maps from the baselines; however, they fail to reconstruct the inputs when the PriPHiT ($\epsilon=1$) method is used. It is interesting to see that in the outputs of the deep reconstruction attacks, the makeup indicators are changed, when PriPHiT is applied.

### v2

@@ D White-Box Reconstruction Attack
E Privacy Budget Analysis
Figs. 16 and 17 show the effect of the privacy budget $\epsilon$ on the accuracy of the cloud analyzer in comparison to the accuracy of the cloud classification attacker for the smiling versus gender and having an open mouth versus wearing makeup experiments on the CelebA dataset, respectively. In the PriPHiT method, reducing the privacy budget decreases the classification attacker’s accuracy on the sensitive content significantly while causing a minimal cost in the analyzer accuracy on the desired content.
F Qualitative Results of the Reconstruction Attack in Mouth Open vs. Wearing Makeup Experiment
Figs. 18 and 19 show examples of the results of the reconstruction attacks together with the unseen inputs from the testing subset for the mouth open versus wearing heavy makeup experiment on the CelebA dataset. We see that the output of both attacks can reconstruct the inputs with makeup indicators using the extracted feature maps from the baselines; however, they fail to reconstruct the inputs when the PriPHiT ($\epsilon=1$) method is used. It is interesting to see that in the outputs of the deep reconstruction attacks, the makeup indicators are changed, when PriPHiT is applied.